OpenStudy (anonymous):

Help will give a medal! What is the equation of the quadratic graph with a focus of (−4, 17/8) and a directrix of y = 15/8? f(x) = −2x^2 + 16x − 24 f(x) = −2x^2 + 15x − 2 f(x) = 2x^2 + 12x − 10 f(x) = 2x^2 + 16x + 34

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@DeltaForce @DecentNabeel @pokamons @AmyRoseRules

OpenStudy (anonymous):

@PeachesP

OpenStudy (anonymous):

@Loser66 ?

OpenStudy (anonymous):

he word directrix is the clue that we are dealing with a parabola. A parabola is a set of points that are equally distant from the focus (a,b) and the directrix d. So for each point (x,y) on the parabola: (x - a)^2 + (y - b)^2 = (y - d)^2 x^2 - 2ax + a^2 + y^2 - 2by + b^2 = y^2 - 2dy + d^2 Subtract y^2 from both sides... x^2 - 2ax + a^2 - 2by + b^2 = -2dy + d^2 Now I would just substitute in the the values (a = -4, b = 17/8, d = 15/8)... x^2 - 2(-4)x + (-4)^2 - 2(17/8)y + (17/8)^2 = -2(15/8)y + (15/8)^2 Simplify... x^2 + 8x + 16 - (17/4)y + 289/64 = -(15/4)y + 225/64 Add (17/4)y to both sides... x^2 + 8x + 16 + 289/64 = (2/4)y + 225/64 x^2 + 8x + 16 + 289/64 = (1/2)y + 225/64 Subtract 225/64 from both sides... x^2 + 8x + 16 + 1 = (1/2)y x^2 + 8x + 17 = (1/2)y Multiply both sides by 2... 2x^2 + 16x + 34 = y ...which is equal to equation D.

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Romnay said it perfectly, its almost exact to what i would say